PLEASE SHOW WITH ALL THE WORKING

Least-Squares method for a linear regression gives following:

Least-Squares method for a linear regression gives following 月 S. where g is the slope of a prediction line +Ax Show that Sn-Σ(x-x)(y--)-J:) I-1
Answer #1

Now

S_{xy} = n s_{xy} = n (\overline{xy} - \overline{x} *\overline{y})

= \sum xy -n * \overline{x} *\overline{y}

= \sum xy -n * \frac{\sum \overline{x}}{n} * \frac{\sum \overline{y}}{n}

= \sum xy - \frac{\sum \overline{x}\sum \overline{y}}{n}